### v1

# On the minimal integral energy of majorants of the Wiener process

Abstract
We consider the asymptotic behavior (over long time intervals) of the minimal integral energy
$|h|_{T}^{\psi}=\int_{0}^{T}\psi(h^{\prime}(t))\,\mathrm{d}t$
of majorants of the Wiener process $W(\cdot)$ satisfying the constraints $h(0)=r$, $h(t)\geq W(t)$ for $0\leq t\leq T$.
The results significantly generalize previous asymptotic estimates obtained for the case of kinetic energy $\psi(u)=u^{2}$, revealing that this case, where the minimal energy grows logarithmically, is a critical one, lying between two different asymptotic regimes.
1 Problem statement and main results
1.1 Introduction
Let $AC[0,T]$ be the space of absolutely continuous functions on the interval $[0,T]$. For an arbitrary convex function $\psi:\mathbb{R}\to\mathbb{R}_{+}$ and any function $h\in AC[0,T]$, consider the energy functional
$|h|_{T}^{\psi}:=\int_{0}^{T}\psi(h^{\prime}(t))\mathrm{d}t.$
The articles [1, 2, 3, 4, 5, 6, 7, 8, 9] investigate the problem of approximating trajectories of various stochastic processes by functions $h(\cdot)$ that minimize a certain type of energy, while satisfying constraints on the proximity between $h(\cdot)$ and the process trajectory. Most often, the approximated process is the Wiener process and the energy considered is kinetic energy, that is, $|h|_{T}^{\psi}$ with $\psi(u)=u^{2}$. The articles [5, 7, 8] study energy-efficient approximations of the Wiener process $W(\cdot)$ under bilateral uniform constraints for arbitrary energy functionals. For $T,r>0$, define the set of admissible approximations
$M^{\pm}_{T,r}:=\left\{h\in AC[0,T]\,\big{|}\,\forall t\in[0,T]:|h(t)-W(t)|\leq r;h(0)=0\right\}$
and the minimal integral energy
$I^{\pm}_{W}(T,r,\psi):=\inf\left\{|h|_{T}^{\psi}\>\big{|}\>h\in M^{\pm}_{T,r}\right\}.$
In [7], it was shown that the approximation at which $\inf\left\{|h|_{T}^{\psi}\>\big{|}\>h\in M^{\pm}_{T,r}\right\}$ is attained—called the taut string—does not depend on the form of $\psi(\cdot)$. The result of [7] was improved in [8], where it was shown that for a fairly wide class of convex functions $\psi(\cdot)$, and for fixed $r>0$, it holds that
$\frac{I^{\pm}_{W}(T,r,\psi)}{T}\stackrel{{\scriptstyle\mathrm{a.s.}}}{{\longrightarrow}}\int_{{\mathbb{R}}}\psi(u)\,\nu(\mathrm{d}u),\quad T\to\infty,$
where $\nu$ is a measure on ${\mathbb{R}}$ that depends on $r$ and admits an explicit expression.
In this paper, we are interested in the behavior of the minimum value of the functional $|\cdot|_{T}^{\psi}$ over approximations of the Wiener process trajectories under fixed unilateral constraints; that is, the set of admissible approximations for given $T,r>0$ takes the form
$M_{T,r}:=\left\{h\in AC[0,T]\,\big{|}\,\forall t\in[0,T]:h(t)\geq W(t)-r;h(0)=0\right\},$
and we are interested in the asymptotic behavior of the minimal integral energy
$I_{W}^{\psi}(T,r):=\inf\left\{|h|_{T}^{\psi}\>\big{|}\>h\in M_{T,r}\right\}$
as $T\to\infty$. The asymptotics of this quantity for $\psi(u)=u^{2}$ is studied in [10].
For technical convenience, we shift the initial value of the approximation to the point $r$, so that the approximating function lies above the trajectory of the process $W(\cdot)$. Define
$M_{T,r}^{\prime}:=\left\{h\in AC[0,T]\,\big{|}\,\forall t\in[0,T]:h(t)\geq W(t);h(0)=r\right\}.$
Since the sets $M_{T,r}$ and $M_{T,r}^{\prime}$ differ only by a constant shift, it is clear that
$I_{W}^{\psi}(T,r)=\inf\left\{|h|_{T}^{\psi}\>\big{|}\>h\in M_{T,r}^{\prime}\right\}.$
We will show below that, unlike in the bilateral constraint case, the value of $r$ does not affect the asymptotic behavior of $I_{W}^{\psi}(\cdot,r)$.
If the minimum of $\psi(\cdot)$ is not attained at zero or is not equal to 0, then $|\cdot|_{T}^{\psi}$ is minimized by a function that is eventually a non-random linear function. Therefore, only functions $\psi(\cdot)$ satisfying $\psi(0)=0$ and attaining their minimum at zero are of interest.
It turns out that the behavior of the minimal energy is determined by the values of the function $\psi(\cdot)$ in a right neighborhood of zero. As follows from Theorem 9 below, if for small $u>0$ the values of $\psi(u)$ are significantly smaller than $u^{2}$, then the minimal integral energy is finite over the entire time axis; that is, $\sup_{T>0}I_{W}^{\psi}(T,r)<\infty$ almost surely. For this reason, the main focus is on convex functions $\psi(\cdot)$ that are $\varkappa$-regularly varying at zero, with $\varkappa\in[1,2]$; that is, functions of the form
$\psi(u)=u^{\varkappa}\,\theta(u),$ (1)
where $\theta(\cdot)$ is slowly varying at zero. The behavior of $I_{W}^{\psi}(T,r)$ primarily depends on the value of the parameter $\varkappa$, with the cases $\varkappa=2$ and $\varkappa<2$ being fundamentally different. To express the conditions imposed on $\psi(\cdot)$, it is convenient to use the Karamata representation for the function $\theta(\cdot)$ (see [11, Theorem 1.2]); namely, for $u\in(0,1)$,
$\theta(u)=c(u)\exp\left(\int_{u}^{1}\frac{\varepsilon(v)}{v}\,\mathrm{d}v\right),$ (2)
where $c(\cdot)$ is a positive bounded measurable function $c(\cdot)$ with a nonzero limit at zero, and $\varepsilon(\cdot)$ is a bounded measurable function with zero limit at zero.
In subsection 1.2, we state results for the case $\varkappa\in[1,2)$. In subsection 1.3, we state results for the case $\varkappa\geq 2$.
In Section 2, we establish a connection between unilateral energy-efficient approximations of arbitrary continuous functions and their minimal concave majorants. We also derive key properties of the minimal concave majorant of the Wiener process, relying heavily on the results of Groeneboom [12]. Section 3 contains the proofs of the theorems for the case $\varkappa\in[1,2)$. Section 4 contains the proofs of the theorems for the case $\varkappa\geq 2$.
1.2 Results for $\varkappa\in[1,2)$
The first theorem describes lower bounds for the lower limit of $I_{W}^{\psi}(T,r)$ as $T\to\infty$; specifically, it identifies deterministic functions of $T$ that are almost surely eventually exceeded by $I_{W}^{\psi}(T,r)$. In other words, any approximation of $W(\cdot)$ almost surely requires more energy to remain above $W(\cdot)$ over the entire time interval $[0,T]$ for all sufficiently large $T$.
Theorem 1
Let $\psi(\cdot)$ be a convex function of the form (1) with $\varkappa\in[1,2)$, and suppose $\theta(\cdot)$ admits the Karamata representation (2) for $u\in(0,1)$. Assume further that there exists $\alpha\in(0,1)$ such that, for all sufficiently small $v>0$,
$\varepsilon(v)<\frac{1}{|\log v|^{\alpha}}.$ (3)
Let $g(\cdot)$ be a slowly varying function at infinity satisfying, for some $\xi_{\circ}>2-\varkappa>\xi^{\circ}>0$,
$(\log b)^{-\xi_{\circ}}\ll g(b)\ll(\log b)^{-\xi^{\circ}},\quad b\to\infty$ (4)
and
$\int^{\infty}\frac{g(b)^{1/(2-\varkappa)}}{b}\,\mathrm{d}b<\infty.$ (5)
Then, for any $r>0$, it holds that
$\liminf_{T\to\infty}\frac{I_{W}^{\psi}(T,r)}{T\psi(T^{-1/2})g(T)}\stackrel{{\scriptstyle\text{a.s.}}}{{=}}\infty.$
Remark 2
Examples of admissible functions $\theta(\cdot)$ include $|\log(\cdot)|^{\alpha}$ with $\alpha\in\mathbb{R}$ and   $\exp\left(\pm|\log(\cdot)|^{\alpha}\right)$ with $\alpha\in(0,1)$. Note that the constraint on $\varepsilon(\cdot)$ is unilateral, meaning it controls only the potential growth of $\theta(u)$ as $u\downarrow 0$. The theorem imposes no restrictions on the possible decay of $\theta(u)$ in the same limit.
The next theorem describes upper bounds for the lower limit of $I_{W}^{\psi}(T,r)$ as $T\to\infty$; specifically, it identifies deterministic functions of $T$ that almost surely infinitely often exceed $I_{W}^{\psi}(T,r)$. In other words, these functions demonstrate how small the energy cost of approximating $W(\cdot)$ can be over certain long time intervals, while still remaining above $W(\cdot)$.
Theorem 3
Let $\psi(\cdot)$ be a convex function of the form (1) with $\varkappa\in[1,2)$, and suppose $\theta(\cdot)$ admits the Karamata representation (2) for $u\in(0,1)$. Assume further that there exists $\alpha\in(0,1)$ such that, for all sufficiently small $v>0$,
$|\varepsilon(v)|<\frac{1}{|\log v|^{\alpha}}.$
Let $g(\cdot)$ be a slowly varying function at infinity satisfying, for some $\xi_{\circ}>2-\varkappa>\xi^{\circ}>0$,
$(\log b)^{-\xi_{\circ}}\ll g(b)\ll(\log b)^{-\xi^{\circ}},\quad b\to\infty$ (6)
and
$\int^{\infty}\frac{g(b)^{1/(2-\varkappa)}}{b}\,\mathrm{d}b=\infty.$ (7)
Then, for any $r>0$, it holds that
$\liminf_{T\to\infty}\frac{I_{W}^{\psi}(T,r)}{T\psi(T^{-1/2})g(T)}\stackrel{{\scriptstyle\text{a.s.}}}{{=}}0.$
Remark 4
Examples of admissible functions $\theta(\cdot)$ include $|\log(\cdot)|^{\alpha}$ with $\alpha\in\mathbb{R}$ and   $\exp\left(\pm|\log(\cdot)|^{\alpha}\right)$ with $\alpha\in(0,1)$. In contrast to Theorem 1, the constraint on $\varepsilon(\cdot)$ in this case is bilateral.
In the case $\psi(u)=u$, Theorems 1 and 3 together yield the well-known Hirsch integral test [13, 14].
Corrolary 5
For any slowly varying function $g(\cdot)$ at infinity satisfying condition (6) with $\varkappa=1$, it holds that
$\liminf_{T\to\infty}\frac{\max_{0\leq t\leq T}W(t)}{T^{1/2}g(T)}\stackrel{{\scriptstyle\text{a.s.}}}{{=}}\begin{cases}0,&\text{if }\int^{\infty}\frac{g(b)}{b}\,\mathrm{d}b=\infty,\\ \infty,&\text{if }\int^{\infty}\frac{g(b)}{b}\,\mathrm{d}b<\infty.\end{cases}$
We now describe the behavior of the upper limit of $I_{W}^{\psi}(T,r)$ as $T\to\infty$. The following theorem provides maximal necessary energy values required for approximating $W(\cdot)$ from some sufficiently large time onward, and characterizes energy levels that may be insufficient.
Theorem 6
Let $\psi(\cdot)$ be a convex function of the form (1) with $\varkappa\in[1,2)$, and suppose $\theta(\cdot)$ admits the Karamata representation (2) for $u\in(0,1)$. Assume further that there exists $\alpha\in(0,1)$ such that, for all sufficiently small $v>0$,
$\varepsilon(v)>-\frac{1}{|\log v|^{\alpha}}.$
Then, for any $r>0$, it holds almost surely that
$\frac{2^{-\varkappa/2}}{1-\varkappa/2}\geq\limsup\limits_{T\to\infty}\frac{I_{W}^{\psi}(T,r)}{T\psi\left(\sqrt{\frac{\log\log T}{T}}\right)}\geq 2^{\varkappa/2}.$
Remark 7
Note that the constraint on $\varepsilon(\cdot)$ is unilateral, regulating only the possible decay of $\theta(u)$ as $u\downarrow 0$. The theorem imposes no restrictions on the possible growth of $\theta(\cdot)$ in the same limit.
Example 8
For $\psi(u)=u^{\varkappa}$ with $\varkappa\in[1,2)$ and any $r>0$, it holds almost surely that
$\frac{2^{-\varkappa/2}}{1-\varkappa/2}\geq\limsup\limits_{T\to\infty}\frac{I_{W}^{\psi}(T,r)}{T^{1-\varkappa/2}(\log\log T)^{\varkappa/2}}\geq 2^{\varkappa/2}.$
1.3 Results for $\varkappa\geq 2$
In contrast to the case $\varkappa\in[1,2)$, when $\varkappa\geq 2$, the minimal energy $I_{W}^{\psi}(T,r)$ is almost surely either bounded on the entire positive real axis or exhibits an unbounded deterministic asymptotic growth. The key quantity in this alternative is
$m_{\psi}(b):=\int^{b}_{1}a\psi(a^{-1})\mathrm{d}a,\quad b\geq 1.$
The conditions that determine the behavior of $I_{W}^{\psi}(T,r)$ are
$\lim_{b\to\infty}m_{\psi}(b)<\infty,$ (8)
$\lim_{b\to\infty}m_{\psi}(b)=\infty.$ (9)
Due to the monotonicity of $m_{\psi}(\cdot)$, exactly one of these conditions holds.
Finite energy case
It turns out that the boundedness of $m_{\psi}(\cdot)$ implies the boundedness of $I_{W}^{\psi}(\cdot,r)$.
Theorem 9
Let $r>0$, and let $\psi(\cdot)$ be a convex function attaining its minimal value 0 at zero. If condition (8) holds, then, almost surely, the random variable $I_{W}^{\psi}(T,r)$ is uniformly bounded in $T$ and increases monotonically to its finite limit as $T\to\infty$.
Example 10
The functions $\psi(u)=u^{\varkappa}$ with $\varkappa>2$, as well as $\psi(u)=u^{2}|\log u|^{\alpha}$ with $\alpha<-1$, satisfy condition (8), along with functions $\psi(\cdot)$ that decay to 0 faster as $u\downarrow 0$.
Exact asymptotics case
In the case where $m_{\psi}(\cdot)$ is unbounded, the exact asymptotic growth of $I_{W}^{\psi}(\cdot,r)$ can be determined.
Theorem 11
Let $\psi(\cdot)$ be a convex function of the form (1) with $\varkappa=2$ and some function $\theta(\cdot)$, which is not necessarily slowly varying. Suppose that for some $\alpha^{\diamond},\alpha_{\diamond}$, one of the following conditions holds:
$-\frac{1}{2}<\alpha^{\diamond},\quad\alpha_{\diamond}\leq\alpha^{\diamond}<\alpha_{\diamond}+\frac{1}{3},$ (10)
or
$-1<\alpha^{\diamond}\leq-\frac{1}{2},\quad\alpha_{\diamond}\leq\alpha^{\diamond}<3\alpha_{\diamond}+2.$ (11)
Assume that
$|\log u|^{\alpha_{\diamond}}\ll\theta(u)\ll|\log u|^{\alpha^{\diamond}},\quad u\to 0.$ (12)
If condition (9) is satisfied, then for any $r>0$, it holds that
$\lim\limits_{T\to\infty}\frac{I_{W}^{\psi}(T,r)}{m_{\psi}\left(T^{1/2}\right)}\stackrel{{\scriptstyle\text{a.s.}}}{{=}}1\,.$
Example 12
The functions $\psi(u)=u^{2}|\log u|^{\alpha}\,\Upsilon(|\log u|)$ with $\alpha>-1$ and some slowly varying function $\Upsilon(\cdot)$ at infinity satisfy condition (9). For $\Upsilon(\cdot)\equiv 1$, by Theorem 11, for any $r>0$, it holds that
$\lim_{T\to\infty}\frac{I_{W}^{\psi}(T,r)}{\left(\log T\right)^{1+\alpha}}\stackrel{{\scriptstyle\text{a.s.}}}{{=}}\frac{1}{(1+\alpha)2^{1+\alpha}}.$
The special case $\alpha=0$ and $\Upsilon(\cdot)\equiv 1$ (kinetic energy) is considered in [10].
2 Preliminaries
2.1 Concave majorants as energy-efficient approximations
The energy-efficient approximation under a unilateral constraint can be characterized in terms of the minimal concave majorant (MCM) of the approximated function and does not depend on $\psi(\cdot)$. Let $w:[0,T]\mapsto{\mathbb{R}}$ be a continuous function. Then the corresponding MCM $\overline{w}(\cdot)$ is the smallest concave function satisfying the condition
$\overline{w}(t)\geq w(t),\qquad 0\leq t\leq T.$
Proposition 13
Let $r>w(0)$ and let $\psi(\cdot)$ be a convex function attaining its minimal value 0 at zero. Assume that, for each $\delta>0$, $\overline{w}(\cdot)$ is a piecewise linear function with a finite number of non-decreasing segments on $[\delta,T]$. Then the problem $|h|_{T}^{\psi}\to\min$ under the constraints $h(0)=r$ and
$h(t)\geq w(t),\qquad 0\leq t\leq T,$
admits a solution $\chi^{\text{ }}_{*}(\cdot)$ of the following form.
(a) If $r\geq\max_{0\leq t\leq T}w(t)$, then $\chi^{\text{ }}_{*}(t)\equiv r$.
(b) If $r<\max_{0\leq t\leq T}w(t)$, then $\chi^{\text{ }}_{*}(\cdot)$ is defined differently on three segments. On the initial segment, $\chi^{\text{ }}_{*}(\cdot)$ is an affine function whose graph passes through the point $(0,r)$ and is tangent to the graph of $\overline{w}(\cdot)$. Then $\chi^{\text{ }}_{*}(\cdot)$ coincides with $\overline{w}(\cdot)$ until the first time the maximum of $w(\cdot)$ is attained. Finally, after this point, $\chi^{\text{ }}_{*}(\cdot)$ remains constant.
Moreover, if $\psi(\cdot)$ is strictly convex, then this solution is unique.
Proof of the proposition: Case (a) is trivial, so consider case (b). Suppose there exists ${\chi^{\text{ }}_{\circ}}(\cdot)$ such that $|\chi^{\text{ }}_{\circ}|_{T}^{\psi}<|\chi^{\text{ }}_{*}|_{T}^{\psi}$. Define the function
$\chi(t):=r+\int_{0}^{t}\vec{\chi}^{\text{ }\prime}_{\circ}(\vartheta)\mathrm{d}\vartheta,\qquad 0\leq t\leq T,$
where $\vec{\chi}^{\text{ }\prime}_{\circ}(\cdot)$ is the decreasing rearrangement of the function $\max\{\chi^{\prime}_{\circ}(\cdot),0\}$. By construction, $\chi(\cdot)$ is a concave non-decreasing function, with $\chi(0)=r$ and $\chi(t)\geq\chi^{\text{ }}_{\circ}(t)\geq w(t)$ for all $t\in[0,T]$. Therefore, $\chi(\cdot)$ satisfies the problem’s constraints. On the other hand,
$|\chi|_{T}^{\psi}=\int_{0}^{T}\psi(\vec{\chi}^{\text{ }\prime}_{\circ})\mathrm{d}t=\int_{0}^{T}\psi(\max\{\chi_{\circ}^{\prime}(t),0\})\mathrm{d}t\leq|\chi^{\text{ }}_{\circ}|_{T}^{\psi}<|\chi^{\text{ }}_{*}|_{T}^{\psi}.$
Since $\chi^{\text{ }}_{*}(\cdot)$ is the minimal convex non-decreasing function satisfying the constraints of the problem, we have
$\chi(t)\geq\chi^{\text{ }}_{*}(t),\qquad 0\leq t\leq T.$
Let $k_{*}$ be the number of affine components composing $\chi^{\text{ }}_{*}(\cdot)$, finite by assumption. Let $\{f_{k}(\cdot)\}_{k=1}^{k_{*}}$ be affine functions on $[0,T]$ such that
$\chi^{\text{ }}_{*}(t)=\min_{1\leq k\leq k_{*}}\{f_{k}(t)\},\qquad 0\leq t\leq T.$
For $k_{*}\geq k\geq 1$ and $t\in[0,T]$, denote $\chi^{\text{ }}_{0}(t):=\chi(t)$ and
$\chi^{\text{ }}_{k}(t):=\min\left\{\min_{1\leq i\leq k}\{f_{i}(t)\},\chi(t)\right\}.$
For $k_{*}>k\geq 0$, we have
$\chi^{\text{ }}_{k+1}(t)=\min\left\{\chi^{\text{ }}_{k}(t),f_{k+1}(t)\right\},\qquad 0\leq t\leq T.$
We will prove that for $k_{*}>k\geq 0$, it is true that
$|\chi^{\text{ }}_{k+1}|_{T}^{\psi}\leq|\chi^{\text{ }}_{k}|_{T}^{\psi}.$
Let $[t_{1},t_{2}]\subset[0,T]$ be the maximal interval on which $\chi^{\text{ }}_{k}(\cdot)\geq f_{k+1}(\cdot)$. Due to the convexity of $\chi^{\text{ }}_{k}(\cdot)$, this interval is well-defined.
Using Jensen’s inequality, we obtain
$(t_{2}-t_{1})\int_{t_{1}}^{t_{2}}\frac{1}{t_{2}-t_{1}}\psi(\chi^{\prime}_{k}(t))\mathrm{d}t\geq(t_{2}-t_{1})\psi\left(\int_{t_{1}}^{t_{2}}\frac{1}{t_{2}-t_{1}}\chi^{\prime}_{k}(t)\mathrm{d}t\right)$
$=$ $(t_{2}-t_{1})\psi\left(\frac{\chi^{\text{ }}_{k}(t_{2})-\chi^{\text{ }}_{k}(t_{1})}{t_{2}-t_{1}}\right)\geq(t_{2}-t_{1})\psi\left(\frac{f_{k+1}(t_{2})-f_{k+1}(t_{1})}{t_{2}-t_{1}}\right)$
$=$ $\int_{t_{1}}^{t_{2}}\psi(f_{k+1}^{\prime}(t))\mathrm{d}t.$
Thus, we have
$\int_{t_{1}}^{t_{2}}\psi(\chi^{\prime}_{k}(t))\mathrm{d}t\geq\int_{t_{1}}^{t_{2}}\psi(f_{k+1}^{\prime}(t))\mathrm{d}t.$
It follows that $\chi^{\text{ }}_{k+1}(\cdot)$ satisfies the problem’s constraints and $|\chi^{\text{ }}_{k+1}|_{T}^{\psi}\leq|\chi^{\text{ }}_{k}|_{T}^{\psi}$. Hence, we obtain that $|\chi|_{T}^{\psi}\geq|\chi^{\text{ }}_{k_{*}}|_{T}^{\psi}=|\chi^{\text{ }}_{*}|_{T}^{\psi}$, which leads to a contradiction.
If $\psi(\cdot)$ is strictly convex, then the energy functional $|\cdot|_{T}^{\psi}$ is also strictly convex, ensuring that $\chi^{\text{ }}_{*}(\cdot)$ is the unique minimizer of $|\cdot|_{T}^{\psi}$. $\square$
2.2 Simplified representation of the minimal integral energy
For any function $f:{\mathbb{R}}\to{\mathbb{R}}$ and any set $F\subset{\mathbb{R}}$, define
$\operatorname*{arg\,max}_{t\in F}f(t):=\sup\left\{t_{0}\in F\,\big{|}\,f(t_{0})=\sup_{t\in F}f(t)\right\}.$
Consider the stochastic process